It asks to solve this problem analytical and show graph
x+2y=8
2x+2y=6

Found 2 solutions by mananth, MathTherapy:
Answer by mananth(16946)


x+2y=8
multiply all terms by 2
we get
2x+4y =16
2x+2y=6
subtract the equations
2y = 10
y=5
plug y in x +2y =8
x+10 =8
x= -2
(-2,5) is the solution of the equation
.

Answer by MathTherapy(10552)


It asks to solve this problem analytical and show graph
x+2y=8
2x+2y=6
Pay NO ATTENTION to that other person. It's beyond STUPID to alter any of the 2 equations. 
x + 2y = 8 ------ eq (i)
2x + 2y = 6 ----- eq (ii)
You simply need to SUBTRACT eq (i) from eq (ii) to get: x = - 2.
Substitute - 2 for x in (i) and solve for y to get its value,
That's all there is to it!